Diffusion Morning Report 04/25/11

Ted Berg tberg at stanford.edu
Mon Apr 25 06:12:41 PDT 2011


Good Morning All,
     Looking good this morning. PH3 for Epi passed leak check and will 
be turned on this morning. Tylan LTO had O2 mfs alarms. Thermco 2 had no 
TLC flow. No other major issues at this time.

					
	Date: 	4/25/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*5.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*1.94/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*.62/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*11.62/30* 	** 	**
Tylan 11-LTO 	*yellow* 	*O2 flow issues* 	*17.06/30* 	** 	*2*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	*
* 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*yellow* 	*PH3 back on today* 	** 	** 	*1*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 429 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110425/12086bbe/attachment.html>


More information about the diffusion mailing list